# Training Machine Learning models at the Edge: A Survey

### v2

@@ V Overview of Edge Learning techniques
collaborative learning (V-A), text depth=2ex, text width=27ex, minimum height=1cm [Federated Learning (V-A1)] [Split Learning (V-A2)] [Swarm Learning (V-A3)] [Gossip Learning (V-A3)] ] [Adaptive and fine-tuning
@@
based technique (V-B), text depth=2ex, text width=27ex, minimum height=1cm) [Transfer Learning (V-B1)] [Incremental Learning (V-B2)] [Meta Learning (V-B3)] ] [Model compression (V-C), text width=27ex, minimum height=1cm [Quantization (V-C2)] [Knowledge Distillation (V-C1)] [Model Pruning (V-C3)] ] [Other optimization
@@
methods (V-D), text depth=2ex, text width=27ex, minimum height=1cm [Binary Neural Networks (V-D1)] [Spiking Neural Networks (V-D2)] [Forward-Forward Algorithm (V-D3)] ] ]
@@
Figure 3: An overview of Edge Learning techniques
V-A Distributed and Collaborative Techniques
@@
In this section, we will explore distributed techniques to train machine learning models at the edge. They work by leveraging the computational capabilities of multiple edge devices, and aggregating their results, instead of relying on a single resource constrained device.
V-A1 Federated Learning
@@
Federated learning (FL), offers a transformative approach to decentralized model training. In the context of edge learning, where data is distributed across numerous edge devices, FL enables collaborative training without centralizing sensitive data [37]. This technique involves training a shared model across these devices by iteratively updating it based on local data, while preserving data privacy and minimizing communication overhead. FL is used extensively for training ML models at the edge [9], and has been adopted for multiple use-cases that range from detecting cyberattacks [38, 39] to spam detection [40, 41], or used in smart cities [42, 43] and autonomous vehicles [44]. Overall, privacy preservation is usually the main objective for FL [45, 46, 47].
@@
In order to train an FL algorithm, an aggregation method is needed, Federated Learning aims to generate a global model by aggregating local models from multiple clients, this process combines the individual models to create a generalized model that represents the collective knowledge of all clients. The main two aggregation methods being, FedSGD (Federated Stochastic Gradient Descent) and FedAVG (Federated Averaging) [37]. However, other approaches have been proposed over the years such as EdgeFed [48] that reduce the number of calculations and time consumption of FedAvg by separating the process of updating the local model that is supposed to be completed independently by mobile devices. FedSel [49] an alternative to FedSGD that aims to solve its dimension dependency problem, by selecting Top-k dimensions according to their contributions in each iteration of federated SGD. Other approaches include MTFeeL [50], FedDynamic [51], FedNets [52], FedCom [53], FedGPO [54], and FedOVA [55].
@@
Despite its growing popularity and multiple benefits, traditional FL models suffer from some limitations, for instance Non-IID (non-independent and identically distributed) data, often negatively impacts the performance of the global model [51]. FL is also susceptible to malicious and low-quality users [56], emerging new classes with completely unseen data distributions from any edge device, whose data cannot be accessed by the global server or other users [57] as well as single node failure [58, 59], channel bandwidth bottlenecks [60], and scaling issues for increasing network size [58]. To solve the low performance with Non-IID data challenge, Hybrid Federated Learning approaches (Hybrid FL) have been proposed [61], where very small amounts of data is shared from the participants. Other approaches that aim to solve this problem include FedNets [52], FedDynamic [51] and [62] that proposes a one-shot Neural Architecture Search technique. Pairwise Correlated Agreement (PCA) [56], on the other hand, is a method that aims to evaluate individual users’ contribution to avoid malicious and low-quality contributions from users. Sharma et. al [63] proposes a framework to study different noise patterns in user feedback, and explore noise-robust mitigation techniques for training FL models, and [57] propose a unified zero-shot framework to handle emerging classes in edge devices.
Over-The-Air Federated Learning (OTA-FL) [64] is another concept in FL that optimizes wireless channels for simultaneous edge device transmission of model updates, aggregating the models over the air in a ”one-time” manner, and minimizing resource consumption, this approach enhances communication efficiency through shared spectral resources and achieve efficient and low-latency global model aggregation [65]. Multiple approaches using OTA-FL have been developed in the past years. Mital and Gunduz [66] proposes an over-the-air computation (OAC) scheme for block-fading MACs that modulate the source to fit the available channel bandwidth in a wideband channel. [67] evaluates the system performance of FEEL with misaligned OAC. [68] propose a Hierarchical Over-The-Air Federated Learning (HOTAFL) method to mitigate the low performances of OTA-FL by the presence of mobile users located far away from the server. Other OTA-FL approaches include [69, 70, 71, 72].
@@
Hierarchical Federated Learning (HFL) and Hierarchical Federated Edge Learning (HFEL) are an extension of FL that introduces a multi-level architecture [73]. These architectures allow for more efficient communication and computation trade-offs [74], faster model training, and reduced energy consumption by offloading tasks to edge servers for partial model aggregation to reduce network traffic [75]. [76] propose a H-FEEL training algorithms that address challenges in helper scheduling and communication resource allocation, and [75] propose a task offloading approach based on data and resource heterogeneity to improve training performance and reduce system cost. Other variations of FL/FEEL include Blind Federated Edge Learning [77], Modular Federated Learning [78], and Clustered Federated Learning (CFL) [79] that will be presented in more details in section VI-A.
@@
Training language and multimedia models at the edge using FL is increasingly explored. While the training of Large Language Models (LLMs) using FL is still experimental, some approaches have been proposed such as FATE-LLM [80] and FwdLLM [81] that aim to fine-tune billion-sized language models across mobile devices using FL. On the other hand, there has been more work on the training of smaller language models such as BERT [82] with FL, such as FedBERT [83] that uses FL and SL approaches for pre-training BERT in a federated way, while FedSplitBERT [84] handles heterogeneous data and decreases the communication cost by splitting BERT encoder layers into a local part trained locally by the client and a global part trained by aggregating gradients of multiple clients. On the other hand, FedSPAM [40], fine tune a distilBERT model [85] using FL on mobile devices to detect spams in SMSes. In computer vision, FedVKD [86] was proposed as a federated knowledge distillation training algorithm to train small CNN models on edge devices and periodically transfers their knowledge to a large server-side vision transformer encoder via knowledge distillation, and [87] introduces an FL approach for visual classification with real-world data distribution. Finally, training audio models at the edge using FL is wildly explored for tasks such as speech recognition [88, 89, 90, 91, 92] or audio classification [57, 93].
@@
Federated Learning is also used alongside other techniques presented in this survey, such as split learning [94], meta learning [95, 96], transfer learning [97], knowledge distillation [25, 98, 99], and Quantization [100, 99, 84], etc.
V-A2 Split Learning
@@
Split Learning (SL) is an alternative collaborative learning method that enables clients to train ML models without sharing their data. In contrast to FL, where models are trained on local data from different devices then aggregated on a central server, SL divides the model into sections. Each section is trained on a different client. Instead of transferring raw data, only the weights of the last layer of each section are sent to the next client. This process ensures model improvement while maintaining better data and model privacy than FL due to the model architecture split between clients and the server, the split also makes SL a better option for resource-constrained environments, at the cost of slower processing than FL due to its relay-based training [94].
@@
Split Learning at the edge has been increasingly studied in the last couple of years (see figure 5), such studies include, SplitEasy [101] a framework for training ML models on mobile devices using SL, [102] that proposes a data protection approach for SL without compromising the model accuracy, and [103] proposes a split edge learning based method by designing an online model splitting and resource provisioning game scheme that aims to minimize the total time cost of participating devices. Adaptive SL (AdaSplit), is a branch of SL that aims to solve its shortcomings compared to FL (slower processing, suboptimal performance and high bandwidth consumption), AdaSplit addresses these challenges by eliminating the transmission of gradients from the server to the client, resulting in a smaller payload and reduced communication cost, and allowing the client to update only sparse partitions of the server model, adapting to the variable resource budgets of different clients which decreases the computation cost and improves performance across heterogeneous clients [104, 105]. Other AdaSplit approaches for edge learning include ARES [106] and [107]. Finally, SL has been combined with FL multiple times in order to eliminate both techniques’ inherent drawbacks, notably in [94, 108, 109, 110, 111, 86, 112].
V-A3 Other Collaborative Learning methods
@@
Several distributed learning techniques have been proposed as alternatives to FL and SL. Swarm learning, an innovative approach integrating artificial and biological intelligence, addresses challenges in distributed ML for the edge. This method efficiently utilizes signal processing and communication techniques to operate in real-time within large-scale edge IoT environments, offering advantages in overcoming communication bottlenecks, diverse data, non-convex optimization, and privacy concerns [113]. Another approach to swarm learning, CB-DSL [114], a Communication-efficient and Byzantine-robust Distributed Swarm Learning technique, was introduced to deal with Non-IID data issues and byzantine attacks. Another noteworthy distributed learning method is gossip learning, which, like other collaborative methods, doesn’t require transferring data outside edge devices. However, unlike FL and other methods, gossip learning operates without a central server for model aggregation and lacks reliance on central control [115]. Notable extensions to gossip learning, such as the one proposed by [116], enhance the algorithm by incorporating additional memory for storing local caches of model updates, making it more suitable for mobile devices.
V-B Adaptive and Fine-tuning based Techniques
@@
In this section, we discuss techniques for efficiently adapting and fine-tuning pre-trained ML models at the edge without requiring complete retraining. The focus of these techniques is on preserving privacy and achieving personalized performance, while reducing computational overhead by keeping the heaviest part of the training in the cloud or edge servers, either using public datasets or ethically collected data. We explore three approaches: transfer learning, incremental learning, and meta learning. These methods enable edge devices to leverage previously acquired knowledge, adapt to local data distributions, and the continuous improvement of the models on the edge.
V-B1 Transfer Learning
@@
Transfer Learning (TL) is an ML technique where knowledge gained from solving one problem is applied to a different, yet related, problem. Instead of building models from scratch, TL employs pre-trained models on large datasets to extract valuable insights, such as learned features or representations. These insights are then used to enhance the performance of a new task, especially when limited data is available for that task. By capitalizing on existing knowledge, TL accelerates model training, improves generalization, and proves exceptionally useful in domains where data scarcity poses a challenge [117].
@@
In the context of edge learning, TL is a prominent technique used to fine-tune ML models based on local data in an edge device, and can serve as a fully on-device alternative to collaborative learning methods such as FL or SL that distribute the training across different devices [31]. Here are some state-of-the-art methods for transfer learning techniques for edge learning. TinyTL (Tiny-Transfer-Learning) [30] aims to achieve on-device learning memory efficiency, which is especially important for low memory edge devices. It achieves that by freezing the weights of the model and only learning a memory-efficient bias module, thus removing the need to store the intermediate activations. Similarly, RepNet [118] proposes an intermediate feature re-programming of a pre-trained model with a tiny Reprogramming Network to develop memory-efficient on-device transfer learning, and MobileTL [119] proposes a memory and computationally efficient on-device transfer learning method for models built with Inverted Residual Blocks. [120] propose an edge CNN framework for 5G industrial edge networks, with the CNN model trained in advance in an edge server, which is further fine-tuned based on the limited datasets uploaded from the devices with the aid of transfer learning. And [121] proposes a runtime convergence monitor to achieve massive computational savings in the practical on-device training workloads. Multiple approaches also focus on combining TL and FL, to create federated transfer learning algorithms [122], that aim to leverage FL for privacy preservation, and use transfer learning to train a well-performing local model despite users usually having not enough data for that by training the base model with a public dataset and passing it to the federated users to be fine-tuned for the target task [97, 123]. Finally, [124] proposes freeze and reconfigure, a TL method for on-device training of a BERT model.
V-B2 Incremental Learning
@@
Incremental learning also referred to as continual learning or life-long learning, involves continuously updating and expanding a model’s knowledge as new data becomes available. Unlike traditional batch learning, where models are trained from scratch on entire datasets, incremental learning dynamically incorporates new information without discarding previously acquired knowledge [125] [126], and can be used to reduce/overcome the well-known issue of catastrophic forgetting in deep neural networks [127, 128, 129].
@@
There has been considerable attempts of implementing incremental learning in the context of edge learning. These include: learning with sharing [130] which aims to reduce the training complexity and memory requirements while achieving high accuracy during the incremental learning process and bypass the considerable memory requirements that can make incremental learning unsuited for edge devices; PILOTE [127] that trains an incremental learning model on edge devices for human activity recognition; [131] introduces an incremental algorithm based on transfer learning and k-nearest neighbor to support the on-device learning; RIANN [132] is an indexing and search system for graph-based approximate nearest neighbor algorithm for mobile devices; and RILOD [133] which aims to incrementally train an existing object detection model to detect new object classes without losing its capability to detect old classes, to avoid catastrophic forgetting. RILOD distils three types of knowledge from the old model to mimic the old model’s behaviour on object classification, bounding box regression and feature extraction, and it was implemented under both edge-cloud and edge-only setups [133]. There are a variety of promising approaches and directions for incremental learning on the edge from combining it with other techniques (such as Federated learning [134, 135], meta learning [136, 135] and compression methods [137, 138, 139]) to Sparse [140] or Distributed Continual Learning [138].
V-B3 Meta-learning

### v3

@@ III Overview of Edge Learning techniques
collaborative learning (III-A), text depth=2ex, text width=27ex, minimum height=1cm [Federated Learning (III-A1)] [Split Learning (III-A2)] [Swarm Learning (III-A3)] [Gossip Learning (III-A3)] ] [Adaptive and fine-tuning
@@
based technique (III-B), text depth=2ex, text width=27ex, minimum height=1cm) [Transfer Learning (III-B1)] [Incremental Learning (III-B2)] [Meta Learning (III-B3)] ] [Model compression (III-C), text width=27ex, minimum height=1cm [Quantization (III-C2)] [Knowledge Distillation (III-C1)] [Model Pruning (III-C3)] ] [Other optimization
@@
methods (III-D), text depth=2ex, text width=27ex, minimum height=1cm [Binary Neural Networks (III-D1)] [Spiking Neural Networks (III-D2)] [Forward-Forward Algorithm (III-D3)] ] ]
@@
Figure 2: A taxonomy of techniques used to enable and/or optimize edge learning
III-A Distributed and Collaborative Techniques
@@
In this section, we will explore distributed techniques to train ML models at the edge. They work by leveraging the computational capabilities of multiple edge devices, and aggregating their results, instead of relying on a single resource constrained device.
III-A1 Federated Learning
@@
Federated Learning (FL), offers a transformative approach to decentralized model training. In the context of edge learning, where data is distributed across numerous edge devices, FL enables collaborative training without centralizing sensitive data [50]. This technique involves training a shared model across these devices by iteratively updating it based on local data, with the objective of preserving data privacy [51, 52, 53]. FL has been widely adopted for edge learning [6], with applications in various domains, including cyberattack detection [54, 55], spam detection [56, 57], smart cities [58, 59], and autonomous vehicles [60].
@@
In order to train an FL algorithm, an aggregation method is needed. Federated learning aims to generate a global model by aggregating local models from multiple clients. This process combines individual models to create a generalized one that represents the collective knowledge of all clients. The main two aggregation methods being, Federated Stochastic Gradient Descent (FedSGD) and Federated Averaging (FedAVG) [50]. However, other approaches have been proposed over the years such as EdgeFed [61] which reduces FedAvg’s computational overhead by separating the process of updating the local model that is supposed to be completed independently by mobile devices, or FedSel [62] which addresses FedSGD’s dimension dependency problem, by selecting Top-k dimensions according to their contributions in each iteration. Other approaches include MTFeeL [63], FedDynamic [64], FedNets [65], FedCom [66], FedGPO [67], and FedOVA [68].
@@
Despite its growing popularity and multiple benefits, traditional FL models suffer from some limitations. For instance, Non-IID (Non-Independent and Identically Distributed) data often negatively impacts the performance of the global model [64]. FL is also vulnerable to malicious and low-quality users [69], emerging new classes with completely unseen data distributions whose data cannot be accessed by the global server or other users [70] as well as single node failure [71, 72], channel bandwidth bottlenecks [73], and scaling issues for increasing network size [71]. To solve the low performance with Non-IID data challenge, Hybrid FL approaches have been proposed [74], where very small amounts of data is shared between participants. Other approaches that aim to solve this problem include FedNets [65], FedDynamic [64] and [75] which proposes a one-shot neural architecture search technique. In contrast, pairwise correlated agreement [69], is a method that aims to evaluate individual users’ contribution to avoid malicious and low-quality contributions from users. Sharma et al. [76] proposes a framework to study different noise patterns in user feedback, and explore noise-robust mitigation techniques for training FL models. Finally, [70] proposes a unified zero-shot framework to handle emerging classes in edge devices.
@@
Hierarchical federated learning, an extension of FL, introduces a multi-level architecture [77], which enables more efficient communication and computation trade-offs [78]. Furthermore, they facilitate faster model training and reduce energy consumption by offloading tasks to edge servers for partial model aggregation to reduce network traffic [79]. For instance, [80] proposes a hierarchical training algorithms that address challenges in helper scheduling and communication resource allocation. While  [79] developed a task offloading approach based on data and resource heterogeneity to improve training performance and reduce system cost. Other variations of FL include blind federated edge learning [81], modular federated learning [82], and clustered federated learning [83] which will be presented in more details in section IV-A.
@@
There is a growing interest in training language and multimedia models at the edge using FL, with several approaches being proposed in recent years. While the training of Large Language Models (LLMs) using FL is still experimental, some approaches have been proposed such as FATE-LLM [84] and FwdLLM [85] which aim to fine-tune a billion parameter language models across mobile devices using FL. In contrast, relatively smaller language models like BERT [86] have been extensively explored using FL, with approaches such as FedBERT [87] that uses FL and SL approaches for pre-training BERT in a federated way. FedSplitBERT [88] addresses the challenges of heterogeneous data and decreases the communication cost by splitting BERT encoder layers into two parts. A local part trained on the client-side and a global part trained by aggregating gradients of multiple clients. Another example is FedSPAM [56], which fine-tunes a distilBERT model [89] using FL on mobile devices to detect spams in SMSes. In computer vision, FedVKD [90] was proposed as a federated knowledge distillation training algorithm to train small CNN models on edge devices. Periodically, the knowledge from these models is transferred to a large server-side vision transformer encoder via knowledge distillation. On the other hand, [91] introduces an FL approach for visual classification with real-world data distribution. Finally, training audio models at the edge using FL is wildly explored for tasks such as speech recognition [92, 93, 94, 95, 96] or audio classification [70, 97].
One important concept usually related to FL is differential privacy. Differential privacy is a privacy preservation technique that involves adding artificial noise to protect individual privacy while maintaining model utility [98][99] provides a detailed examination of differential privacy while [100] and [101] provides an exploration of differential privacy in the context of FL.
@@
FL is also used alongside other techniques presented in this survey, such as split learning [102], meta learning [103, 104], transfer learning [105], knowledge distillation [28, 106, 107], and Quantization [108, 107, 88], etc.
III-A2 Split Learning
@@
Split learning offers an alternative approach to collaborative learning. In contrast to FL, which involves training models on local data from different devices and aggregating them on a central server, split learning takes a different approach. Specifically, it divides the model into sections, with each section trained on a different client or server, and instead of transferring raw data, only the weights of the last layer of each section are sent to the next client. This process ensures model improvement while maintaining better data and model privacy than FL, thanks to the model architecture split between clients and the server. Additionally, this split makes split learning a more suitable option for resource-constrained environments, where computational resources are limited. However, this approach comes at the cost of slower processing than FL, which is due to its relay-based training [102].
@@
In recent years, there has been a growing interest in split learning at the edge, as evident from the increasing number of studies in this area (see figure 3). For instance, SplitEasy [109] is a framework that enables the training ML models on mobile devices using split learning. Another paper, [110] proposes a data protection approach for split learning without compromising the model accuracy. Additionally, [111] proposes an online model splitting method with resource provisioning game scheme which aims to minimize the total time cost of participating devices. Adaptive split learning, is a branch of split learning that aims to overcome its shortcomings compared to FL. Specifically, it addresses these challenges by eliminating the transmission of gradients from the server to the client, resulting in a smaller payload and reduced communication cost, and allowing the client to update only sparse partitions of the server model, adapting to the variable resource budgets of different clients which decreases the computation cost and improves performance across heterogeneous clients [112, 113]. Other adaptative approaches include ARES [114] and [115]. Finally, split learning has been combined with FL in multiple approaches in order to eliminate both techniques’ inherent drawbacks, notably in [102, 116, 117, 118, 119, 90, 120].
III-A3 Other Collaborative Learning methods
@@
Several distributed learning techniques have been proposed as alternatives to FL and split learning. Swarm learning, an innovative approach integrating artificial and biological intelligence, addresses challenges in distributed ML for the edge. This method efficiently utilizes signal processing and communication techniques to operate in real-time within large-scale edge IoT environments, offering advantages in overcoming communication bottlenecks, diverse data, non-convex optimization, and privacy concerns [121]. Another approach to swarm learning, CB-DSL [122], a Communication-efficient and Byzantine-robust Distributed Swarm Learning technique, was introduced to deal with Non-IID data issues and byzantine attacks. Another noteworthy distributed learning method is gossip learning, which, like other collaborative methods, doesn’t require transferring data outside edge devices. However, unlike FL and other methods, gossip learning operates without a central server for model aggregation and lacks reliance on central control [123]. Notable extensions to gossip learning, such as the one proposed by [124], enhance the algorithm by incorporating additional memory for storing local caches of model updates, making it more suitable for mobile devices.
III-B Adaptive and Fine-tuning based Techniques
@@
In this section, we discuss techniques for efficiently adapting and fine-tuning pre-trained ML models at the edge without requiring a complete retraining. The focus of these techniques is on preserving privacy and achieving personalized performance, while reducing computational overhead by keeping the heaviest part of the training in the cloud or edge servers, either using public datasets or ethically collected data. We explore three approaches: transfer learning, incremental learning, and meta learning. These methods enable edge devices to leverage previously acquired knowledge, adapt to local data distributions, and the continuous improvement of the models on the edge.
III-B1 Transfer Learning
@@
Transfer learning is an ML technique where knowledge gained from solving one problem is applied to a different, yet related, problem. Instead of building models from scratch, transfer learning employs pre-trained models on large datasets to extract valuable insights, such as learned features or representations. These insights are then used to enhance the performance of a new task, especially when limited data is available for that task. By capitalizing on existing knowledge, transfer learning accelerates model training, improves generalization, and proves exceptionally useful in domains where data scarcity poses a challenge [125]. For a more detailed understanding of transfer learning, readers are encouraged to review the surveys [126, 127, 125].
@@
In the context of edge learning, transfer learning is a prominent technique used to fine-tune ML models based on local data in an edge device. This approach serves as a fully on-device alternative to collaborative learning methods that distribute the training across different devices [44]. Notable state-of-the-art methods for transfer learning in edge learning include tiny-transfer learning [43], which addresses the critical issue of memory efficiency in low-memory edge devices. This is achieved this by freezing the weights of the model and only learning a memory-efficient bias module, thus removing the need to store the intermediate activations. Similarly, RepNet [128] proposes an intermediate feature re-programming of a pre-trained model with a tiny reprogramming network to develop memory-efficient on-device transfer learning. MobileTL [129] also proposes a memory and computationally efficient on-device transfer learning method, specifically designed for models built with inverted residual blocks. Additionally, [130] propose an edge CNN framework for 5G industrial edge networks, with the CNN model trained in advance in an edge server, which is further fine-tuned based on the limited datasets uploaded from the devices with the aid of transfer learning, and [131] proposes a runtime convergence monitor to achieve massive computational savings in the practical on-device training workloads. Multiple approaches also focus on combining transfer learning with FL, to create federated transfer learning algorithms [132], that aim to leverage FL for privacy preservation, and use transfer learning to train a well-performing local model despite users usually having not enough data for that by training the base model with a public dataset and passing it to the federated users to be fine-tuned for the target task [105, 133]. Finally, [134] proposes freeze and reconfigure, a transfer learning method for on-device training of a BERT model.
III-B2 Incremental Learning
@@
Incremental learning also known as continual learning or life-long learning, involves continuously updating and expanding a model’s knowledge as new data becomes available. Unlike traditional batch learning, where models are trained from scratch on entire datasets, incremental learning dynamically incorporates new information without discarding previously acquired knowledge [135], and can be used to reduce/overcome the well-known issue of catastrophic forgetting in deep neural networks [136, 137, 138]. Readers seeking a more thorough understanding of incremental learning are directed to [139].
@@
There have been considerable attempts of implementing incremental learning in the context of edge learning. These include: learning with sharing [140] which aims to reduce the training complexity and memory requirements while achieving high accuracy during the incremental learning process and bypass the considerable memory requirements that can make incremental learning unsuited for edge devices; PILOTE [136] which trains an incremental learning model on edge devices for human activity recognition; [141] introduces an incremental algorithm based on transfer learning and k-nearest neighbor to support the on-device learning; RIANN [142] is an indexing and search system for graph-based approximate nearest neighbor algorithm for mobile devices; and RILOD [143] which aims to incrementally train an existing object detection model to detect new object classes without losing its capability to detect old classes, to avoid catastrophic forgetting. RILOD distills three types of knowledge from the old model to mimic the old model’s behaviour on object classification, bounding box regression and feature extraction, and it was implemented under both edge-cloud and edge-only setups [143]. There are a variety of promising approaches and directions for incremental learning on the edge from combining it with other techniques (such as FL [144, 145], meta learning [146, 145] and compression methods [147, 148, 149]) to sparse [150] or distributed continual learning [148].
III-B3 Meta-learning